# 3D Beamforming Through Joint Phase-Time Arrays

### v1

@@ Abstract
High-frequency wide-bandwidth cellular communications over mmW and sub-THz offer the opportunity for high data rates, however, it also presents high pathloss, resulting in limited coverage. To mitigate the coverage limitations, high-gain beamforming is essential. Implementation of beamforming involves a large number of antennas, which introduces analog beam constraint, i.e., only one frequency-flat beam is generated per transceiver chain (TRx). Recently introduced joint phase-time array (JPTA) architecture, which utilizes both true time delay (TTD) units and phase shifters (PSs), alleviates analog beam constraint by creating multiple frequency-dependent beams per TRx, for scheduling multiple users at different directions in a frequency-division manner. One class of previous studies offered solutions with “rainbow” beams, which tend to allocate a small bandwidth per beam direction. Another class focused on uniform linear array (ULA) antenna architecture, whose frequency-dependent beams were designed along a single axis of either azimuth or elevation direction. In this paper, we present a novel 3D beamforming codebook design aimed at maximizing beamforming gain to steer radiation toward desired azimuth and elevation directions, as well as across sub-bands partitioned according to scheduled users’ bandwidth requirements. We provide both analytical solutions and iterative algorithms to design the PSs and TTD units for a desired subband beam pattern. Through simulations of the beamforming gain, we observe that our proposed solutions outperform the state-of-the-art solutions reported elsewhere. This work was done in part while O. Yildiz was an intern at Samsung Research America.
@@ I Introduction
Higher frequency mmWave bands are used in current and next-generation wireless networks due to larger bandwidth availability to provide high data rates [1]. However, these frequency bands experience signal degradation due to higher path loss and shadowing effects, which could be resolved by using large antenna arrays with directional beamforming [2].
@@
The disadvantage of directional beamforming is that traditional architectures like phased antenna arrays (PAA) cause high scheduling latency, owing to analog beam constraint, i.e., only one frequency-flat beam is generated per transceiver chain (TRx, or called RF chain). Traditional architectures serve different user directions in a time-division manner. To serve multiple users in one time slot, massive antennas and digital beamforming relying on multiple TRxs can be beneficial but they lead to high power consumption and cost.
@@
We focus on frequency-dependent beamforming architecture, which could be realized by leaky-wave antennas or true-time-delay (TTD) units [3]. Since leaky-wave antennas are cumbersome and inefficient [4], we consider an architecture with TTD units, which is called joint phase time array (JPTA) architecture [5]. This architecture consists of TTD units in addition to the phase shifters (PSs) used in PAA. By using JPTA, beam training can be performed using a single OFDM symbol [6]. JPTA can also eliminate the beam squint/split in wideband communications [4, 7]. In addition, JPTA can extend coverage area and increase cell or user throughput, by offering more per-user scheduling opportunities thanks to multiple frequency-dependent beams [8].
@@
Frequency-dependent beamforming could be beneficial either during the initial access phase or the data communication phase. In [6] and [9], the authors showed that the initial access pilots can be transmitted towards all directions (across azimuth and elevation directions) in one time-domain resource. Jain et al. [10] and Ratnam et al. [5] improve the data communication phase by steering the frequency-dependent 2D beams to multiple users’ angle of arrivals (AoAs). This way, multiple users at different directions are scheduled in a frequency-division-multiplexing (FDM) manner, even when the antenna panel is associated with only one TRx.
Our goal is to design a method to derive a 3D frequency-dependent beamforming codebook, comprising phase and delay parameters using JPTA architecture. Our design is to schedule multiple users according to their bandwidth requirement for the data communication phase. A summary of our contributions is as follows:
@@
  • We analytically derive a closed-form solution for 3D beamforming design by defining the phase and delay values in two different ways: joint and separated according to their interdependency, see Sec. III-A and III-B.
@@
  • We improve the performance beyond analytical solutions, by applying greedy and gradient descent algorithms and we take the analytically derived values as initial values. See Sec. III-C and III-D.
@@
  • We provide 3D beamforming gain simulations to compare the performance of our methods with the 3D extension of the state-of-the-art 2D beamforming codebook design by [5]. We demonstrate that our iterative algorithms outperform the state-of-the-art by providing more resilient results over different scenarios, see Sec. IV.
@@ II System Model
Figure 1: Demonstration of a joint phase-time array, and the joint and separated beam design approaches.
@@
Phase precision, $\beta$ 6 bits
@@
We consider 3D beam design for uplink communication with a system bandwidth ${\rm BW}$ around a center frequency $f_{c}$. There is a single base station (BS) and $N_{u}$ users in the environment. The BS is equipped with a single radio-frequency (RF) chain and a rectangular antenna array, comprising $N_{\rm az}$ antenna elements in the azimuth direction and $N_{\rm el}$ antenna elements in the elevation direction, as shown in Fig. 1. The antenna spacing is half-wavelength at the center frequency $f_{c}$. The antenna array steering vector for $(y,z)$-th antenna element is as follows
$\mathbf{a}_{R}(\theta_{\rm az},\theta_{\rm el},f_{m})$ $=\frac{1}{\sqrt{N_{\rm az}N_{\rm el}}}e^{-j\pi\frac{f_{m}}{f_{c}}(y\sin\theta_{\rm az}\sin\theta_{\rm el}+z\cos\theta_{\rm el})},$ (1)
where $f_{m}$ is the subcarrier frequency, $\theta_{\rm az}$ represents AoA in the azimuth plane, and $\theta_{\rm el}$ denotes the AoA in the elevation plane.
@@ II-A Joint Phase Time Array Architecture
We utilize JPTA architecture to design a frequency-dependent 3D beamforming codebook since this architecture consists of PSs and TTD units, which enables control over both the angle and frequency domains. We assume that every antenna element in the rectangular array has a separate TTD and PS; therefore, different phase and delay values can be assigned to each, see Fig. 1.
@@
$\bm{\omega}=\frac{1}{\sqrt{N_{\rm az}N_{\rm el}}}\bm{\omega}_{\text{phase}}(y,z)\odot\bm{\omega}_{\text{delay}}(y,z,f_{m}),$ (2)
where PS weights are $\bm{\omega}_{\text{phase}}(y,z)=e^{j\phi_{y,z}}$ for a given phase $\phi_{y,z}$ and TTD combiner weights are $\bm{\omega}_{\text{delay}}(y,z,t)=\delta(t-\tau_{y,z})$ in the time domain and are $\bm{\omega}_{\text{delay}}(y,z,f_{m})=e^{-j2\pi f_{m}\tau_{y,z}}$ in the frequency domain for a given delay $\tau_{y,z}$.
@@
$G(\theta_{\rm az},\theta_{\rm el},f_{m})$ $=\frac{1}{N_{\rm az}N_{\rm el}}\lVert\bm{a}_{R}^{H}(\theta_{\rm az},\theta_{\rm el},f_{m})\bm{\omega}(y,z,f_{m})\rVert^{2}$ (3)
$=\frac{1}{N_{\rm az}N_{\rm el}}\left\lVert\sum_{y=0}^{N_{\rm az}-1}\sum_{z=0}^{N_{\rm el}-1}e^{jh_{y,z}-j\Omega(y,z)}\right\rVert^{2},$
with the definition of $h_{y,z}\triangleq\phi_{y,z}+2\pi f_{m}\tau_{y,z}$ and $\Omega(y,z)\triangleq\pi\frac{f_{m}}{f_{c}}(y\sin\theta_{\rm az}\sin\theta_{\rm el}+z\cos\theta_{\rm el})$.
@@ II-B Problem Formulation
We aim to maximize the 3D frequency-dependent beamforming gain of $N_{u}$ users for a given bandwidth allocation of each user by configuring PSs and TTD units. The desired bandwidth ratio of the $i$-th user is denoted as $\alpha_{i}$, which corresponds to $\lfloor\alpha_{i}\times\frac{\rm BW}{\Delta f}\rfloor$ subcarriers where ${\Delta f}$ is the subcarrier spacing. To assess the performance of the beam, we employ the log-mean of the average user gain as a metric, as delineated in
$G_{l}\triangleq\sum_{i=1}^{N_{u}}10\log_{10}G_{{\rm mean},i}$ (4)
for the definition of
$G_{{\rm mean},i}\triangleq\frac{1}{|\mathcal{F}_{m,i}|}\sum_{m\in\mathcal{F}_{m,i}}G(\theta_{{\rm az},i},\theta_{{\rm el},i},f_{m}),$ (5)
where the AoAs of the $i^{\rm th}$ user is $(\theta_{{\rm az},i},\theta_{{\rm el},i})$ and assigned subcarrier set is $\mathcal{F}_{m,i}$. The utilization of logarithms in this context further promotes fairness among various users.
Thus, the problem formulation is
$\max_{\bm{\tau},\phi}G_{l}$ (6)
Here, $\bm{\phi}$ and $\bm{\tau}$ correspond to matrices containing the phase and delay values for each antenna element.
@@
We provide several methods to design the 3D beamforming by solving the problem defined in (6). We first outline an analytical derivation to assign the phase and delay values of each antenna element independent of the values of other antenna elements, referred to as joint, in Sec. III-A. Then, we define a similar solution when the phase and delay values in each row and column of antenna elements are correlated, specified as separated, in Sec. III-B. Consequently, we provide two iterative optimization methods to improve upon the analytical derivations, denoted as the greedy algorithm in Sec. III-C and the gradient descent algorithm, in Sec. III-D.
@@ III-A Joint Analytical Solution
We devise a solution for the problem outlined in (6) by establishing a set of linear equations corresponding to each antenna element, similar to [10], given that the phase and delay of antenna elements are set independently.
It is of significance to recognize that the maximum beamforming gain, referred to (3), would be obtained if the equality of $h_{y,z}=\Omega(y,z)$ holds for all subcarriers, where $y$ and $z$ are integers in $[0,N_{\rm az}-1]$ and $[0,N_{\rm el}-1]$, respectively. For analytical convenience, we assume that there are $M+1$ subcarriers and we assign the subcarriers in range $[f_{\alpha_{i-1}(M+1)},f_{\alpha_{i}(M+1)})$ to $i$-th user. Therefore, we formulate a system of linear equations that encompasses an equation for each subcarrier given by $\mathbf{A}\mathbf{x}_{y,z}=\mathbf{b}_{y,z}$, with $\mathbf{x}_{y,z}\triangleq[\phi_{y,z},2\pi\Delta f\tau_{y,z}]^{T}$, and
$\mathbf{A}\triangleq\begin{bmatrix}1&1&\cdots&1\\ -M/2&-M/2+1&\cdots&M/2\\ \end{bmatrix}^{T}.$ (7)
Under the assumption of $f_{m}/f_{c}\approx 1$, $\mathbf{b}_{y,z}$ is defined as,
$\mathbf{b}_{y,z}\triangleq[\mathbf{b}^{T}_{y,z,1},\mathbf{b}^{T}_{y,z,2},\cdots,\mathbf{b}^{T}_{y,z,N_{u}}]^{T},$ (8)
@@
where $\nu_{y,z,i}\triangleq y\pi\sin\theta_{{\rm az},i}\sin\theta_{{\rm el},i}+z\pi\cos\theta_{{\rm el},i}$ is the antenna array steering value for the $i$-th user and
$k_{y,z,i}=\begin{cases}0&i=1\\ k_{y,z,i-1}+{round}\left(\frac{\nu_{y,z,i-1}-\nu_{y,z,i}}{2\pi}\right)&{\rm else,}\end{cases}$ (10)
@@
Given that above defined linear system of equations leads to an overdetermined linear system of equations, finding a unique solution to satisfy every equation is not possible. Thus, we introduce the error term, denoted as $\mathbf{e}_{y,z}\triangleq\mathbf{A}\mathbf{x}_{y,z}-\mathbf{b}_{y,z}$. To maximize the beamforming gain, we need to minimize the error $\mathbf{e}_{y,z}$ for all combinations of $y$ and $z$. To fulfill this, we explore two approaches by employing different norm values for error quantification such as the square norm and the infinity norm.
@@ III-A1 Least Squares Solution
We redefine the problem formulation as $\min_{\bm{\tau},\bm{\phi}}\lVert\mathbf{e}_{i}(\bm{\theta}_{\rm az},\bm{\theta}_{\rm el},\mathcal{F}_{m})\rVert_{2}$ and by definition, the closed form solution is $\widehat{x_{y,z}}=(\mathbf{A}^{T}\mathbf{A})^{-1}\mathbf{A}^{T}\mathbf{b}_{y,z}$, which leads to the phase values of
$\phi_{y,z}=\sum_{i=1}^{N_{u}}\alpha_{i}(\nu_{y,zi,i}+2\pi k_{y,z,i}),$ (11)
@@
$M_{\rm sum}(i,\bm{\alpha})=\sum_{m=0}^{\sum_{l=1}^{i}\alpha_{l}M-M/2}m-\sum_{m=0}^{\sum_{l=1}^{i-1}\alpha_{l}M-M/2}m.$ (13)

### v3

@@ Abstract
High-frequency wideband cellular communications over mmWave and sub-THz offer the opportunity for high data rates. However, it also presents high path loss, resulting in limited coverage. High-gain beamforming from the antenna array is essential to mitigate the coverage limitations. The conventional phased antenna arrays (PAA) cause high scheduling latency owing to analog beam constraints, i.e., only one frequency-flat beam is generated. Recently introduced joint phase-time array (JPTA) architecture, which utilizes both true-time-delay (TTD) units and phase shifters (PSs), alleviates analog beam constraints by creating multiple frequency-dependent beams for scheduling multiple users at different directions in a frequency-division manner. One class of previous studies offered solutions with “rainbow” beams, which tend to allocate a small bandwidth per beam direction. Another class focused on uniform linear array (ULA) antenna architecture, whose frequency-dependent beams were designed along a single axis of either azimuth or elevation direction. This paper presents a novel 3D beamforming design that maximizes beamforming gain toward desired azimuth and elevation directions and across sub-bands partitioned according to scheduled users’ bandwidth requirements. We provide analytical solutions and iterative algorithms to design the PSs and TTD units for a desired subband beam pattern. Through simulations of the beamforming gain, we observe that our proposed solutions outperform the state-of-the-art solutions reported elsewhere. This work was done in part while O. Yildiz was an intern at Samsung Research America.
@@ I Introduction
Higher frequency mmWave bands are used in current and next-generation wireless networks due to larger bandwidth (BW) availability to provide high data rates [1]. However, these frequency bands experience signal degradation due to higher path loss and shadowing effects, which could be resolved using large antenna arrays with directional beamforming [2].
@@
The disadvantage of directional beamforming is that traditional architectures like phased antenna arrays (PAA) cause high scheduling latency owing to analog beam constraints, i.e., only one frequency-flat beam is generated per radio-frequency (RF) chain. Traditional architectures serve different user directions in a time-division-multiplexing (TDM) manner. To serve multiple users in one timeslot, digital beamforming relying on multiple RF chains can be beneficial, but they lead to high power consumption and cost.
@@
We focus on frequency-dependent beamforming architecture, which could be realized by leaky-wave antennas or true-time-delay (TTD) units [3]. Since leaky-wave antennas are cumbersome and inefficient [4], we consider an architecture with TTD units, which is called joint phase time array (JPTA) architecture [5, 6]. This architecture consists of TTD units and phase shifters (PSs). Using JPTA, beam training can be performed using a single OFDM symbol [7]. JPTA can also eliminate the beam squint/split in wideband communications [4, 8]. In addition, JPTA can extend coverage area and increase cell or user throughput by offering more per-user scheduling opportunities thanks to multiple frequency-dependent beams [6].
@@
Frequency-dependent beamforming could be beneficial during both the initial access and data communication phases. In [7] and [9], the authors showed that the initial access pilots could be transmitted towards all directions (across azimuth and elevation directions) in one time-domain resource. In [10] and [5], the data communication phase is improved by steering the frequency-dependent 2D beams to multiple users’ angle of arrivals (AoAs). This way, multiple users in different directions are scheduled in a frequency-division-multiplexing (FDM) manner, even when the antenna panel is associated with only one RF chain.
In this work, we propose methods to derive a 3D frequency-dependent beam by optimizing the phase and delay parameters in the JPTA architecture. The formed beam, which is used for data communications, considers the UE locations and their BW requirements, i.e., UEs with higher BW requirements observe high beam gain over a larger BW than UEs with smaller desired BW. A summary of our contributions is as follows:
@@
  • We analytically derive a closed-form solution for 3D beamforming design by defining the phase and delay values in two different ways: joint and separated according to their interdependence, see Sec. III-A and III-B.
@@
  • We improve the performance beyond analytical solutions by applying greedy and gradient descent algorithms which take the analytically derived values as initial values. See Sec. III-C and III-D.
@@
  • We provide 3D beamforming gain simulations to compare the performance of our methods with the 3D extension of the state-of-the-art 2D beamforming codebook design in [5]. We demonstrate that our iterative algorithms outperform the state-of-the-art by providing more resilient results over different scenarios, see Sec. IV.
@@ II System Model
Figure 1: Demonstration of a joint phase-time array and the joint and separated beam design approaches.
@@
Phase shifter precision, $\beta$ 6 bits
@@
We consider 3D beam design for uplink communication with a system bandwidth ${\rm BW}$ around a center frequency $f_{c}$. A single base station (BS) and $N_{u}$ users are in the environment. The BS is equipped with a single RF chain and a rectangular antenna array, comprising $N_{\rm az}$ antenna elements in the azimuth direction and $N_{\rm el}$ antenna elements in the elevation direction, as shown in Fig. 1. The antenna spacing is half-wavelength at the center frequency $f_{c}$. The antenna array steering value for $(y,z)$-th antenna element is as follows
$\left[\bm{a}_{R}(\theta_{\rm az},\theta_{\rm el},f_{m})\right]_{y,z}$ $=e^{j\pi\frac{f_{m}}{f_{c}}(y\sin\theta_{\rm az}\sin\theta_{\rm el}+z\cos\theta_{\rm el})},$ (1)
where $f_{m}$ is the subcarrier frequency, $\theta_{\rm az}$ and $\theta_{\rm el}$ denote the AoA in the azimuth and elevation plane. Here, $[\cdot]_{y,z}$ represents the $(y,z)$-th element of the matrix.
@@ II-A Joint Phase Time Array Architecture
The JPTA architecture is shown in Fig. 1, where every antenna element in the rectangular array has a separate TTD and PS, enabling control over the angle and frequency domains.
@@
$\left[\bm{\omega}(f_{m},\bm{\phi},\bm{\tau})\right]_{y,z}=\frac{1}{\sqrt{N_{\rm az}N_{\rm el}}}e^{j(\phi_{y,z}+2\pi f_{m}\tau_{y,z})},$ (2)
where $\bm{\tau}$ $(\bm{\phi})$ corresponds to a matrix containing each antenna element’s delay (phase) value. The PS weights are $e^{j\phi_{y,z}}$ for a given phase $\phi_{y,z}$ and TTD combiner weights are $e^{j2\pi f_{m}\tau_{y,z}}$ for a given delay $\tau_{y,z}$.
@@
$G(\theta_{\rm az},\theta_{\rm el},f_{m},\bm{\phi},\bm{\tau})$ (3)
$=$ $\lVert\bm{a}_{R}^{H}(\theta_{\rm az},\theta_{\rm el},f_{m})\bm{\omega}(f_{m},\bm{\phi},\bm{\tau})\rVert^{2}$
$=$ $\frac{1}{N_{\rm az}N_{\rm el}}\left\lVert\sum_{y=0}^{N_{\rm az}-1}\sum_{z=0}^{N_{\rm el}-1}e^{jh(f_{m},\phi_{y,z},\tau_{y,z})-j\Omega(y,z,f_{m},\theta_{\rm az},\theta_{\rm el})}\right\rVert^{2},$
with the definition of $h(f_{m},\phi_{y,z},\tau_{y,z})\triangleq\phi_{y,z}+2\pi f_{m}\tau_{y,z}$ and $\Omega(y,z,f_{m},\theta_{\rm az},\theta_{\rm el})\triangleq\pi\frac{f_{m}}{f_{c}}(y\sin\theta_{\rm az}\sin\theta_{\rm el}+z\cos\theta_{\rm el})$ for simplification.
@@ II-B Problem Formulation
Recall that our goal is to maximize the 3D frequency-dependent beamforming gain of $N_{u}$ users for a given BW allocation of each user by configuring the PSs and the TTD units. The desired BW ratio of the $i$-th user is denoted as $\alpha_{i}$, which corresponds to $\lfloor\alpha_{i}\times\frac{\rm BW}{\Delta f}\rfloor$ subcarriers where ${\Delta f}$ is the subcarrier spacing. Denote the assigned subcarrier set to $i^{\rm th}$ user as $\mathcal{F}_{m,i}$, and the AoAs as $(\theta_{{\rm az},i},\theta_{{\rm el},i})$. The average beamforming gain of $i^{\rm th}$ user over the assigned subcarrier set $\mathcal{F}_{m,i}$ at the desired direction $(\theta_{{\rm az},i},\theta_{{\rm el},i})$ is thus
$G_{{\rm mean},i}\triangleq\frac{1}{|\mathcal{F}_{m,i}|}\sum_{m\in\mathcal{F}_{m,i}}G(\theta_{{\rm az},i},\theta_{{\rm el},i},f_{m}).$ (4)
To assess the performance of the beam, we use the log-mean of the average user gain as a metric,
$G_{l}\triangleq\sum_{i=1}^{N_{u}}10\log_{10}G_{{\rm mean},i}.$ (5)
The utilization of logarithms in this context further promotes fairness among various users [11]. In addition, the formulation approximates maximizing the total sum rate across the $N_{u}$ users at high SNR since the data rate increases logarithmically with SNR. Finally, the problem formulation is to maximize the above metric over the delay and phase shifter values:
$\left\{\bm{\tau}^{*},\bm{\phi}^{*}\right\}=\operatorname*{arg\,max}_{\bm{\tau},\bm{\phi}}G_{l}.$ (6)
@@
We provide several methods to design the 3D beam by solving the problem defined in (6). First, we outline an analytical method to jointly determine the phase and delay values in azimuth and elevation directions, referred to as joint, in Sec. III-A. Next, we define a similar analytical solution when the phase and delay values in the azimuth and elevation directions are designed separately, referred to as separated, in Sec. III-B. Consequently, we propose two iterative optimization techniques to enhance the analytical solutions: denoted as the greedy algorithm in Sec. III-C and the gradient descent algorithm, in Sec. III-D.
@@ III-A Joint Analytical Solution
In this section, we devise a solution for the problem outlined in (6) by establishing a set of linear equations corresponding to each antenna element, similar to [10], given that the phase and delay of the antenna elements are set independently.
Looking back at (3), we can see that the maximum beamforming gain for a given direction $(\theta_{\rm az},\theta_{\rm el})$ and subcarrier $(f_{m})$ is achieved when $h(\cdot)=\Omega(\cdot)$ for all $y$ and $z$ in $[0,N_{\rm az}-1]$ and $[0,N_{\rm el}-1]$, respectively. Hence, if there is a solution satisfying this equality for each UE over its assigned subcarriers, the maximum beamforming gain is achieved for all UEs.
For analytical convenience, we assume that there are $M+1$ subcarriers and we assign the subcarriers in range $[f_{\alpha_{i-1}(M+1)},f_{\alpha_{i}(M+1)})$ to $i$-th user. Therefore, we formulate a system of linear equations that encompasses an equation for each subcarrier given by $\bm{A}\bm{x}_{y,z}=\bm{b}_{y,z}$, with $\bm{x}_{y,z}\triangleq[\phi_{y,z},2\pi\Delta f\tau_{y,z}]^{T}$, and
$\bm{A}\triangleq\begin{bmatrix}1&1&\cdots&1\\ -M/2&-M/2+1&\cdots&M/2\\ \end{bmatrix}^{T}.$ (7)
Under the assumption of $f_{m}/f_{c}\approx 1$, $\bm{b}_{y,z}$ is defined as,
$\bm{b}_{y,z}\triangleq[\bm{b}^{T}_{y,z,1},\bm{b}^{T}_{y,z,2},\cdots,\bm{b}^{T}_{y,z,N_{u}}]^{T},$ (8)
@@
where $\nu(y,z,i)\triangleq y\pi\sin\theta_{{\rm az},i}\sin\theta_{{\rm el},i}+z\pi\cos\theta_{{\rm el},i}$ is the antenna array steering value for the $i$-th user and
$\begin{split}&k(y,z,i)\\ &\triangleq\begin{cases}0&i=1\\ k(y,z,i-1)+{round}\left(\frac{\nu(y,z,i-1)-\nu(y,z,i)}{2\pi}\right)&{\rm else,}\end{cases}\end{split}$ (10)
@@
The defined linear system of equations leads to an over-determined linear system of equations, so finding a unique solution to satisfy every equation is impossible. Thus, we introduce the error term, denoted as $\bm{e}_{y,z}\triangleq\bm{A}\bm{x}_{y,z}-\bm{b}_{y,z}$. To maximize the beamforming gain, we need to minimize the error $\bm{e}_{y,z}$ for all combinations of $y$ and $z$. To fulfill this, we explore two approaches by employing different norm values for error quantification, such as the square and infinity norms.
@@ III-A1 Least Squares Solution
We redefine the problem formulation as $\min_{\bm{\tau},\bm{\phi}}\lVert\bm{e}_{i}(\bm{\theta}_{\rm az},\bm{\theta}_{\rm el},\mathcal{F}_{m})\rVert_{2}$ and by definition, the closed form solution is $\widehat{\bm{x}_{y,z}}=(\bm{A}^{T}\bm{A})^{-1}\bm{A}^{T}\bm{b}_{y,z}$, which leads to the phase values of
$\phi_{y,z}=\sum_{i=1}^{N_{u}}\alpha_{i}(\nu(y,z,i)+2\pi k(y,z,i)),$ (11)
@@
$M_{\rm sum}(i,\bm{\alpha})\triangleq\sum_{m=0}^{\sum_{l=1}^{i}\alpha_{l}M-M/2}m-\sum_{m=0}^{\sum_{l=1}^{i-1}\alpha_{l}M-M/2}m.$ (13)